OpenStudy (anonymous):

Integrate (x^2-4)/x * dx

OpenStudy (anonymous):

Write \[ \frac{x^2+4}x=x+\frac 4 x \] It is easy now to integrate

OpenStudy (anonymous):

The integral is \[ \frac{x^2}2 +4 \ln|x| +C \]

OpenStudy (anonymous):

Change the + 4ln|x| to - 4ln|x| I misread the problem.

OpenStudy (anonymous):

\[ \int \frac{x^2-4}x dx\int x -4 \frac{1}x dx=\frac {x^2}2 - 4\ln|x| + C \]